OpenStudy (anonymous):

Greetings! Derrivate of sin(2x+pi) Help is much appreciated ! :)

Vocaloid (vocaloid):

the general format for a sin derivative is d(sinx)/dx = cos(x) * d(x)/d(x) so d(sin(2x+pi)/dx = cos(2x+pi)*d(2x+pi) = cos(2x+pi)*2

Vocaloid (vocaloid):

here's an easier way to remember/explain it to find the derivative of sin(a), change the sin to cos, keep the a, then multiply by the derivative of a

OpenStudy (rizags):

wait. think about it... what happens to the graph of sine when you add π

OpenStudy (rizags):

\[\sin(2x+π) = -\sin(2x)\]

OpenStudy (anonymous):

ofc that first equation makes sense, however ur last one, I get but its hard to come to that conclusion, I understand pi jumps 180degrees making it negative, but its not obvious.

OpenStudy (rizags):

therefore we have: \[\large -(\frac{d}{dx}(\sin(2x)))\]

OpenStudy (rizags):

chain rule it^^

OpenStudy (anonymous):

u'v+uv'?

OpenStudy (rizags):

right, like so: \[\frac{d}{dx}(\sin(2x))=\frac{dsin(u)}{du}\frac{du}{dx}\] Im going to let \[u=2x\] can you solve now?

OpenStudy (anonymous):

cos(u)*2 Then u=2x so gives Cos2x*2. U forgot the pi tho, what happened to it?

OpenStudy (rizags):

as i said before, the π disappears when i rewrite the entire thing as -sin(2x)

OpenStudy (rizags):

and make sure you add the negative from before. The final answer is: \[\huge -2\cos(2x)\]

OpenStudy (anonymous):

ah okey, I will remember that, so my first step should always be to get rid of pi? then move on the the chain rule and so forth:) Alright, thank you :)

OpenStudy (rizags):

no prob. Medal please!

OpenStudy (rizags):

yea π is an issue

OpenStudy (anonymous):

Haha there ya go, Always fun to learn when things go this smoothly, u explain well:) Cheers!